OpenStudy (anonymous):

There are 6 girls and 7 boys in a class. A team of 10 players is to be selected from the class. What is the probability that a randomly chosen team has 3 or 4 girls?

OpenStudy (anonymous):

nCr=combination of n things taken r at a time= n!/((n-r)!*r!) 4C2 * 6C3 =6*20 =120 groups If a group is chosen randomly from the 10 people, find the probability that the group will consist exactly of 2 girls and 3 boys.

OpenStudy (anonymous):

@anthonygatesjr

OpenStudy (anonymous):

@latracia thanks

OpenStudy (anonymous):

no problem